By state contract between the Government and the University 240 master students should get education. Number of students is defined by equation Q=K^(1/4)*L(1/4) , where K is the size of the University buildings, L is the number of class hours.
(a) Rent rate is equal to 2, class hour costs 8. Find the lowest costs of education for one student, 240 students.
(b) The Government changes state contract. The University chooses number of students and the Government pays p for each student. Assume that 4% of students do not defend a graduation thesis and the University do not get money for such students. Find p, which creates incentives to educate 240 students.
1
Expert's answer
2019-08-12T09:47:33-0400

a. 240="2^{(\frac{1}{4})}"× 8"(\frac{1}{4})"

240=2.38, cost for 1 student will be = 0.0099

b. 4% × 240= 9.6= 9 students

p= (240-9)× 0.0099= 2.2869